### v1

# xFinder: Robust and Pinpoint Answer Extraction for Large Language Models

Abstract
The continuous advancement of large language models (LLMs) has brought increasing attention to the critical issue of developing fair and reliable methods for evaluating their performance. Particularly, the emergence of subjective or non-subjective cheating phenomena, such as test set leakage and prompt format overfitting, poses significant challenges to the reliable evaluation of LLMs. Since evaluation frameworks often utilize Regular Expression (RegEx) for answer extraction, some models may adjust their responses to comply with specific formats that are easily extractable by RegEx. Nevertheless, the key answer extraction module based on RegEx frequently suffers from extraction errors. This paper conducts a comprehensive analysis of the entire LLM evaluation chain, demonstrating that optimizing the key answer extraction module can improve extraction accuracy, reduce LLMs’ reliance on specific answer formats, and enhance the reliability of LLM evaluation. To address these issues, we propose xFinder, a model specifically designed for key answer extraction. As part of this process, we create a specialized dataset, the Key Answer Finder (KAF) dataset, to ensure effective model training and evaluation. Through generalization testing and evaluation in real-world scenarios, the results demonstrate that the smallest xFinder model with only 500 million parameters achieves an average answer extraction accuracy of 93.42%. In contrast, RegEx accuracy in the best evaluation framework is 74.38%. xFinder exhibits stronger robustness and higher accuracy compared to existing evaluation frameworks. All resources for xFinder are available at https://github.com/IAAR-Shanghai/xFinder.
1 Introduction
Large language models (LLMs) have seen rapid advancements in recent years [1, 2]. To explore the strengths, weaknesses, and differences of various LLMs, numerous researchers have introduced a variety of evaluation benchmarks aimed at evaluating the performance of LLMs across various dimensions [3, 4, 5, 6]. Furthermore, to facilitate comparisons of different LLMs on these benchmarks, several institutions have developed unified LLM evaluation frameworks, such as LM Eval Harness [^1] LM Eval Harness is the backend for Hugging Face’s popular Open LLM Leaderboard [7]. [8] and OpenCompass [9]. These frameworks are capable of uniformly evaluating LLM in reasoning, text generation, and domain-specific knowledge.
Despite the widespread application of the aforementioned evaluation benchmarks and frameworks, some researchers have raised concerns about the fairness and reliability of current LLM evaluation efforts [10, 11]. For instance, the benchmarks may suffer from data contamination and leakage in their test sets [12, 13]. Furthermore, a systematic review has revealed multiple unreliabilities at various stages of the existing evaluation pipeline, as illustrated in Figure 1.
Figure 1: Typical LLM Evaluation Pipeline.
Data Item Transformation
Given the uncertainty of LLM output content, evaluation frameworks often transform datasets to achieve uniform evaluation across different tasks. For example, they may convert datasets into the Multiple-Choice Question-Answer (MCQA) with alphabet options. However, it has been shown that for some tasks, the evaluation results under this MCQA form with alphabet options may differ from those under the open-ended setting  [14].
Question Prompting and LLM Answering
Various institutions may employ different configurations (such as temperature settings, decoding strategies, etc.) and prompts when evaluating LLMs. LLMs are highly sensitive to these settings [12]. Even minor modifications (such as changing the case of letters in options or adding extra spaces) can significantly alter the model’s responses [14, 15, 16, 17, 18], thus affecting the fairness of the evaluation outcomes. Moreover, while employing consistent prompts may ensure a degree of fairness, uniform prompts may not suit all LLMs, potentially contradicting the initial intent of evaluating the models’ inherent capabilities.
Key Answer Extraction and Matching
Current evaluation frameworks predominantly rely on Regular Expression (RegEx) to extract answers from model outputs. For example, they may select content following phrases like "The answer is" as the key response. However, many LLMs fail to generate standardized answers [19], which complicates the extraction of key answers from their outputs. Figure 2 illustrates instances where current evaluation frameworks failed to extract responses correctly. There are two primary scenarios where regular rules fail to extract answers: first, when the model does not produce a relevant response to the evaluation question, and second, when the model’s response does not conform to standards. In the first scenario, it is unreasonable to classify a model’s failure to provide a relevant response as an error, as the evaluation aims to measure the model’s knowledge comprehension and reasoning capabilities, not merely its following ability. In the second scenario, although the LLM’s response effectively addresses the question, the response’s non-standard format impedes correct extraction using RegEx, resulting in ’false judgments’.
Figure 2: Cases where LM Eval Harness and OpenCompass fail in extracting key answers. A/T/C/M stands for tasks with alphabet / short text / categorical label / math options, respectively.
Consequently, existing LLM evaluation practices still face numerous deficiencies in terms of fairness and reliability. How to conduct reliable evaluations has become an urgent issue in the field of LLMs. For extracting and matching key answers, the most effective and reliable methods involve using advanced LLMs (such as GPT-4) alongside manual evaluation [20]. However, due to time and financial constraints, these methods are seldom applied in practical evaluations. As a result, some initiatives have developed models, known as Judge models, through fine-tuning strategies to automatically determine the correctness of LLM responses [21, 22]. Unfortunately, these methods significantly lag behind advanced LLMs or manual evaluation in terms of generalizability.
To address the conflict between fairness and cost, we propose xFinder, a novel evaluation component designed to replace the answer extraction phase of existing frameworks. We have optimized the logic for matching extraction results, thereby enhancing the reliability of evaluations. The core of xFinder involves constructing a large-scale LLM response evaluation dataset and proposing a high-precision, low-parameter model based on this dataset to serve as the evaluator. By replacing the existing RegEx-based answer evaluation methods, xFinder can evaluate LLM outputs more accurately and efficiently, further enhancing the reliability of LLM evaluations.
To construct xFinder, we initially created the Key Answer Finder (KAF) Dataset. The training set is used to fine-tune xFinder, and we successfully fine-tuned 19 LLMs of varying sizes, all of which performed well on both the test and generalization sets. For instance, on the testing set, xFinder-qwen1505 achieved an average extraction accuracy of 96.88%. To ensure the generalizability of xFinder, we constructed a generalization set using outputs from various models on different tasks beyond the original KAF dataset. xFinder-qwen1505 achieved an average extraction accuracy of 93.42% on this dataset, demonstrating the model’s substantial generalizability. Finally, to ensure consistent experimental settings, we utilized xFinder and the traditional evaluation framework’s RegEx answer extraction module to assess 10 LLMs on a common set of evaluation tasks under real-world conditions, subsequently ranking the 10 LLMs’ performance. We found that the RegEx-based key answer extraction components in existing frameworks often fail to extract answers correctly or at all. Moreover, rankings produced by different evaluation frameworks show significant inconsistencies. In contrast, our fine-tuned xFinder models of various sizes achieved high consistency in LLM rankings, demonstrating greater stability in the evaluation results.
We summarize our primary contributions as follows:
  • We provide a comprehensive review of LLM evaluation processes in the industry, identifying critical factors that can lead to unreliable evaluation results.
  • We introduce xFinder, a model specifically designed for key answer extraction. The KAF dataset supports its effective training and evaluation.
  • In our extensive experiments, we demonstrate that RegEx-based evaluation methods are unreliable, while our xFinder model significantly improves reliability.
2 Related Work
From the perspective of LLM capabilities, current evaluations can be categorized into knowledge and capability evaluations [23], vertical domain evaluations [24], and alignment and safety evaluations [25, 26]. When considering the types of questions in benchmark datasets, evaluations can be divided into objective questions (such as multiple-choice, classification, mathematical questions, fill-in-the-blank, etc.) and subjective questions (such as summarization, text continuation, open-ended questions, and code generation). For objective questions, metrics like accuracy and F1-score are widely used. These metrics rely on the ability to exactly match the LLM’s output with the correct answer.
Currently, many popular unified LLM evaluation frameworks exist, including LM Eval Harness [8], OpenCompass [9], OpenAI Evals [27], and UltraEval [28]. When using objective question datasets to evaluate LLMs, these frameworks first conduct data item transformation, converting classification and fill-in-the-blank questions into multiple-choice questions. They then extract and match key answers using RegEx methods. This process either involves literally extracting the option letters for text response evaluation or extracting the probabilities of different option letters during generation for first-token evaluation. However, both methods have notable drawbacks when evaluating objective questions. In text response evaluation, inaccurately extracting the key answer from the LLM’s response can lead to incorrect evaluation results. On the other hand, first-token evaluation may not accurately reflect the LLM’s intended answer, as it overlooks the LLM’s reasoning process [29].
In addition to RegEx-based answer extraction, recent research has increasingly focused on fine-tuning LLMs for evaluation, known as LM-based methods. Specifically, these methods include training evaluation models to select the better response to the same question (Pairwise Selection) [21, 22, 30] and training evaluation models to score a given question directly, response, and reference (Pointwise Grading) [31, 30, 32]. However, these methods have significant drawbacks. The current prompt and fine-tuning dataset designs often result in evaluation LLMs with poor generalization, making them difficult to compare with strong LLMs like GPT-4 [33, 20].
Recently, the concept of reliable evaluation has been proposed  [^2] https://sites.google.com/view/real-info2024/overview. This concept aims to more scientifically assess the authenticity of content generated by LLMs and the potential risks to information systems [34]. However, systematic research and efforts to improve reliable evaluation are still lacking. To the best of our knowledge, our work is the first to conduct a systematic study in this area.
3 Problem Definition
The existing methods described in Section 2 aim to align the scores assigned by evaluation models with human-expected scores. In other words, these methods strive to ensure that the evaluation model’s judgments of LLM outputs match human judgments. Since scoring is subjective, unlike the objective determination of correctness, these methods are more suitable for open-ended tasks like text translation and summarization. In contrast, our method aims to align xFinder’s extraction of key answers from LLM outputs with the key answers expected by humans. Fundamentally, this method is more suitable for tasks with deterministic answers such as multiple-choice questions and mathematical problems. We define the key answer extraction task as follows:
Let $\mathcal{D}$ denote an evaluation dataset consisting of triplets: a question, a set of options, and the correct answer, as shown in Equation 1. Here, $q$ denotes the question, $\mathcal{C}$ denotes the set of options, and $a$ denotes the correct answer. Let $\Sigma$ be the set of all tokens, and $\mathcal{P}$ the power set.
$\mathcal{D}\subset\{(\boldsymbol{q},\mathcal{C},\boldsymbol{a})\mid\boldsymbol{q}\in\Sigma^{*},\boldsymbol{a}\in\Sigma^{*},\mathcal{C}\in\mathcal{P}(\Sigma^{*})\}$ (1)
Each data point $(\boldsymbol{q},\mathcal{C},\boldsymbol{a})$ in $\mathcal{D}$ is fed into the LLM, producing an output $\boldsymbol{y}$, as shown in Equation 2:
$\boldsymbol{y}=\text{LLM}(\boldsymbol{q},\mathcal{C},\boldsymbol{a})$ (2)
If a unique substring in $\boldsymbol{y}$ can be identified as one of the three types—Direct, Prompt wrapped, or Converted question wrapped—then this substring is considered the Key Answer Sentence $\boldsymbol{s}$. Intuitively, $\boldsymbol{s}$ is a sentence that contains the final answer to the question $\boldsymbol{q}$. The definitions of these three types and their corresponding sets are as follows:
Direct
[…]
Prompt wrapped
Here, $\boldsymbol{s}$ consists of relevant prompts combined with the final answer, as detailed in Equation 4. The subset $\mathcal{F}\subset\mathcal{P}(\Sigma^{*})$ represents a collection of prompts guiding to the answer, where the placeholder <final answer> indicates the position of the final answer. The expression $\boldsymbol{s}\circ\boldsymbol{f}$ denotes the new sentence formed by replacing the <final answer> token in $\boldsymbol{f}$ with $\boldsymbol{s}$.
Converted question wrapped
[…]
Particularly, when multiple substrings meet the criteria, it is necessary to determine if a Chain of Thought (CoT) process [^3] CoT (Chain of Thought [35]) refers to the process of progressively reasoning through $\boldsymbol{q}$ to arrive at the final result. exists within $\boldsymbol{y}$. If a CoT process is present in $\boldsymbol{y}$, $\boldsymbol{s}$ is defined as the unique substring that meets the criteria and appears after the CoT process in $\boldsymbol{y}$. If no CoT process is present, $\boldsymbol{s}$ does not exist in $\boldsymbol{y}$.
$\mathcal{S}_{1}$ =$=\{\boldsymbol{c},\tau(\boldsymbol{c})\mid\boldsymbol{c}\in\mathcal{C}\}$ (3)
$\mathcal{S}_{2}$ =$=\{\boldsymbol{s}\circ\boldsymbol{f}\mid\boldsymbol{s}\in\mathcal{S}_{1},\boldsymbol{f}\in\mathcal{F}\}$ (4)
$\mathcal{S}_{3}$ =$=\{\boldsymbol{s}\circ\boldsymbol{x}\mid\boldsymbol{s}\in\mathcal{S}_{1},\boldsymbol{x}\in\zeta(\boldsymbol{q})\}$ (5)
The definition of the key answer $\boldsymbol{k}$ is divided into two scenarios:
  1. 1.
    If $\boldsymbol{s}$ does not exist, $\boldsymbol{k}$ is set to [No valid answer];
  2. 2.
    If $\boldsymbol{s}$ exists (as defined in Equation 6), $\boldsymbol{k}$ is the option within the set $\mathcal{C}$ that best corresponds to what is described by $\boldsymbol{s}$.
[…]
4 Methodology
To effectively fine-tune xFinder, a high-quality dataset is essential. To our knowledge, there is currently no dedicated dataset for Key Answer Extraction. In this section, we will detail the implementation process of xFinder, including the construction of the KAF dataset and the training of xFinder, as illustrated in Figure 3. The KAF dataset encompasses a variety of evaluation tasks, including questions, optional answer ranges, LLM responses to the questions, and the extracted key answers. This dataset is divided into three segments: a training set, a test set, and a generalization set, which are used for fine-tuning, testing, and assessing xFinder’s performance, respectively. The training set comprises 26,900 samples, the test set includes 4,961 samples, and the generalization set contains 4,482 samples. Section 4.1 details the process of LLM response collection, Section 4.2 describes our method and workflow for annotating key answers in the LLM-generated content, and Section 4.3 elaborates on our detailed approach to training xFinder.
Figure 3: Schematic of the research framework. The first three stages correspond to Sections 4.1,  4.2, and  4.3 respectively; the final stage illustrates the replacement of RegEx with xFinder in the current evaluation pipeline. The experiments in Section 5.2 demonstrate the efficacy of our approach within this pipeline. Note: The percentages 80.3%, 77.0%, and 2.2% in the center of the figure around ’unreliable evaluation’ indicate results from the Llama3-8B-Instruct model on the GSM8K benchmark using RegEx evaluation via the LM Eval Harness, OpenCompass, and UltraEval frameworks, respectively, while our method achieves a reliable result of 80.2%.
4.1 LLM Response Generation
As shown in the first part of Figure 3, we initially generate question-response pairs using different LLMs across multiple datasets. Currently, the tasks for evaluating LLM performance are quite diverse. We investigated the leading LLM benchmarks and evaluation reports to cover a wide range of popular evaluation tasks. Ultimately, we selected 18 typical evaluation task datasets, with detailed information provided in Appendix A.1. We categorized these tasks into four types, each with similar option formats, which serve as prompts to induce the LLMs to generate varied responses (as shown in the yellow boxes in Figure 2): (1) Alphabet option tasks. By referring to the design of mainstream evaluation rankings, these tasks are formatted as alphabet options, where each task aims to select the single correct option from multiple possibilities (i.e., A, B, C, D, etc.). (2) Short text option tasks. The requirement for the model to correctly map letters to option texts can affect the assessment of the model’s true capabilities [36]. Therefore, we designed tasks with key answer types as short texts in the form of alphabet options, where the goal is to select the single correct short text from multiple short texts. (3) Categorical label tasks. For traditional classification tasks (e.g., sentiment classification, topic classification, entity classification), we retained their original format, using the labels themselves as options. (4) Math option tasks. Given the output types of mathematical tasks, which could be integers, decimals, or even LaTeX code, we categorized them separately.
To further enhance the heterogeneity and richness of the dataset, we have primarily designed eight types of prompt configurations to elicit diverse responses from the LLMs. These configurations include 0-shot(-restrict444-restrict indicates that the prompt ends with the sentence ”End your final answer with ’The answer is ¡answer¿.’” to standardize the LLM’s response format.), 0-shot-cot(-restrict), 5-shot(-restrict), and 5-shot-cot(-restrict). Specific examples of these prompts are provided in Appendix E.1.
For generating responses, it is crucial to capture a wide range of output content, given that LLMs of different series and sizes exhibit varied reasoning processes and may provide different answers to the same question. Consequently, we utilized 10 distinct LLMs for this purpose. Specific details on these models can be found in Appendix A.2.
It is important to note that the KAF dataset’s training and test sets consist of Question-Response Pairs generated by 10 LLMs across 8 evaluation tasks. To ensure the validity and fairness of the generalization evaluation, the Question-Response Pairs used in the generalization set are derived from responses by 8 LLMs on 11 entirely new evaluation tasks.
4.2 Auto Labelling and Human Recheck
The foundation of constructing the KAF dataset is the extraction of the key answer $\boldsymbol{k}$ from the LLM response $\boldsymbol{y}$, which is then used as the label. We employed distinct annotation strategies for the training, test, and generalization sets. For the training set, to improve the efficiency of the annotation process, a semi-automated procedure was adopted. This involved generating two sets of annotations from GPT-4 using varied prompts E.2. Utilizing the Self-Consistency strategy [37], items with discrepancies between the two annotation rounds, along with all mathematics questions, were consolidated for manual re-annotation using Label Studio [38]. To ensure the validity and reliability of the test and generalization sets, all labels needed to be accurate, necessitating two rounds of manual annotation for these datasets. Detailed information about the annotation methods can be found in Appendix A.3.
4.3 Training xFinder
According to the definition of key answer extraction outlined in Section 3, our goal is to fine-tune xFinder to precisely locate $\boldsymbol{s}$ within $\boldsymbol{y}$ and accurately extract $\boldsymbol{k}$ from $\boldsymbol{s}$. The training input for xFinder includes the question $\boldsymbol{q}$, the LLM response $\boldsymbol{y}$, the key answer type, and the answer range, while the output is the key answer $\boldsymbol{k}$.
During the training process, we primarily utilized the XTuner [39] tool developed by the InternLM team for fine-tuning. To fully exploit the instruction-following potential of xFinder, we meticulously crafted prompts as demonstrated in Appendix E.3. Additionally, to enhance the generalization capability of xFinder, we implemented data augmentation on the KAF training set, which included:
  • Simulating LLM responses to alphabet option questions. We randomly selected 50% of the alphabet option questions from the KAF training set and altered their answer choices by either adding one or two options or removing one option with equal probability.
  • Enriching the variety of prompt forms in Prompt wrapped key answer sentences. From the LLM responses in the KAF training set that contained wrapped key answer sentences, we extracted 10% and substituted the $\boldsymbol{f}$ with other elements from $\mathcal{F}$, such as replacing "The final answer is A" with "Based on the context of the question, A is the most likely answer."
5 Experiments
In this section, we initially fine-tuned 19 LLMs with parameter sizes ranging from 0.5B to 8B using the KAF training set. We compared their accuracy in extracting key answers on both the KAF test and generalization sets against the mainstream evaluation frameworks using RegEx methods. Subsequently, we evaluated the performance of 10 LLMs on 14 evaluation task datasets using xFinder-qwen1505, xFinder-llama38it, and the current mainstream evaluation frameworks, analyzing differences in evaluation results.
5.1 Extraction Accuracy: xFinder vs. RegEx
Results on KAF test set
We employed the QLoRA method from the XTuner framework [39] [40] for fine-tuning the foundation models, with specific tuning parameters detailed in Appendix B. To assess the efficacy of xFinder, we conducted tests using the KAF test set. We set the current mainstream frameworks—OpenCompass, LM Eval Harness, and UltraEval—with their RegEX extraction modules as baselines. The comparison metric is extraction accuracy, defined as the proportion of correctly extracted key answers relative to the dataset size. Partial results are presented in Table 1; for complete experimental results, refer to Table 6 in Appendix C.1.
alphabet option short text categorical label math Overall
OpenCompass 0.8098 / / 0.7481 0.7937
LM Eval Harness 0.8500 0.7540 0.8208 0.7496 0.8113
UltraEval 0.6057 / / 0.4766 0.5722
xFinder-qwen1505 0.9735 0.9683 0.9805 0.9276 0.9688
xFinder-qwen1518 0.9735 0.9781 0.9755 0.9457 0.9712
xFinder-gemma7 0.9751 0.9805 0.9736 0.9367 0.9704
xFinder-chatglm36base 0.9735 0.9793 0.9730 0.9291 0.9684
xFinder-llama38 0.9698 0.9829 0.9692 0.9276 0.9661
Table 1: Comparison of Extraction Accuracy on KAF Test Set: Mainstream Frameworks vs. xFinder Models. Specifically, xFinder-qwen1505, xFinder-qwen1518, xFinder-gemma7, xFinder-chatglm36base, and xFinder-llama38 are models fine-tuned using the KAF training set on Qwen1.5-0.5B, Qwen1.5-1.8B, Gemma-7B, ChatGLM3-6B-base, and Llama-8B, respectively.
It is evident that the xFinder models, fine-tuned using the KAF training set, are highly effective. Furthermore, despite a 16-fold difference in the parameter sizes of the models, the variation in extraction accuracy does not exceed 0.5%. This implies the high quality of the KAF dataset we constructed.
Results on KAF generalization set
To further explore xFinder’s extraction capabilities on not-seen tasks, we conducted tests using the KAF generalization set. Specific results are shown in Table 2. Only two xFinder models are displayed here for brevity, but the complete experimental results can be found in Table 7 of Appendix C.1. The results indicate that xFinder maintains a high level of extraction accuracy on the generalization set. In contrast, mainstream regular expression extractors perform poorly [^5] Here, the extraction accuracy of OpenCompass, LM Eval Harness, and UltraEval has significantly decreased compared to Table 1. This is because the LLM responses in the KAF test set are generated from restricted prompts, making it easier for regular expressions to extract the key answer; however, the responses in the KAF generalization set are generated from non-restricted prompts., especially in extracting key answers for mathematical problems. Even GPT-4 struggles to extract key answers from complex and varied LLM responses accurately.
alphabet option short text categorical label math Overall $\Delta_{acc}$ $\Delta_{acc}/N$
OpenCompass 0.7750 / / 0.6813 0.7438 / /
LM Eval Harness 0.6594 0.7484 0.8381 0.2094 0.6780 / /
UltraEval 0.5945 / / 0.1781 0.3978 / /
GPT-4 0.6578 0.8046 0.6706 0.6703 0.6957 / /
xFinder-qwen1505 0.9477 0.9335 0.9281 0.9234 0.9342 0.0277 0.0554
xFinder-llama38it 0.9547 0.9428 0.9537 0.9547 0.9518 0.0453 0.0057
Table 2: Comparison of Extraction Accuracy on KAF Generalization Set: Mainstream Frameworks, GPT-4, and xFinder Models. Specifically, xFinder-qwen1505 and xFinder-llama38it represent models fine-tuned with the KAF dataset on Qwen1.5-0.5B and Llama3-8b-instruct, respectively.
5.2 Evaluation in Real-World Scenarios
The experiments conclusively demonstrate that xFinder plays a critical role within evaluation frameworks. Next, we will use xFinder for automated evaluations in real-world evaluation tasks instead of only inspecting the extraction accuracy. To balance evaluation efficiency and quality, we calculated two metrics for all xFinder models on the KAF Generalization Set: $\Delta_{acc}$ and $\Delta_{acc}/N$ (where $N$ is the parameter size of the LLM, measured in billions). Based on these metrics, xFinder-qwen1505 (with the highest $\Delta_{acc}/N$) and xFinder-llama38it (with the highest $\Delta_{acc}$) were selected. Here, $\Delta_{acc}$ represents the difference in extraction accuracy on the KAF Generalization Set between the current xFinder and the lowest extraction accuracy among the 19 xFinder models. For specific results, refer to Table 7 in Appendix C.1.
We evaluated the performance of 10 LLMs across 14 mainstream evaluation tasks using xFinder-qwen1505 and xFinder-llama38it, along with regular extraction methods based on OpenCompass, LM Eval Harness, and UltraEval. The comprehensive scoring results are provided in Appendix C.2.
Setup
Currently, the parameter settings for each evaluation framework differ significantly, including preprocessing methods for task datasets, prompt template configurations, and hyperparameter settings for the models being tested. Therefore, to compare the performance of our xFinder extraction method with RegEx methods on these frameworks, we replicated only the RegEx methods on each framework and ensured consistency in other settings. This primarily included: (1) applying uniform preprocessing to the task datasets; (2) employing a consistent zero-shot approach for all evaluation tasks, with specific prompts detailed in Appendix E.1; and (3) selecting the following LLMs for comprehensive evaluation: Phi-2, Baichuan series, Gemma series, Llama series, InternLM series, and Qwen series. Additionally, to ensure the stability of experimental results and maintain consistency in parameter settings across different LLMs, we set the temperature to 0, top_p to 0.9, and top_k to 5.
Figure 4: Bump charts: Changes in LLM rank over different evaluation frameworks.
Results and Analysis
Figure 4 illustrates the ranking changes of LLMs across four tasks based on evaluation results extracted from different frameworks. Additional chart results can be found in Appendix D. From the analysis, we can derive the following insights:
RQ1: Are existing evaluation frameworks reliable? The graph reveals significant discrepancies in the rankings of LLMs obtained using different evaluation frameworks. For instance, in the CommonsenseQA (alphabet) task, InternLM2-1.8B-Chat ranks first in UltraEval but only fifth in OpenCompass. Such disparities in rankings indicate the unreliability of RegEx methods within current mainstream evaluation frameworks, potentially affecting model evaluation and ranking.
RQ2: Is xFinder reliable? Firstly, by comparing our presented checkpoints, xFinder-qwen1505 and xFinder-llama38it, we observe relatively consistent rankings of LLMs across different base models on the same task. This demonstrates the high stability of xFinder, even on CommonsenseQA datasets that have not been used for training. Furthermore, considering the extraction accuracy of key answers from our previous experiments, xFinder exhibits higher accuracy compared to the RegEx methods in other frameworks. This implies that our evaluation results are both stable and accurate. Therefore, it can be concluded that our model demonstrates higher reliability compared to other frameworks.
(additional) RQ3: Is setting evaluation tasks as alphabet options reliable? Existing evaluation frameworks often aim to convert all questions into multiple-choice format with alphabet options. However, as shown in Figure 4, there are significant differences in the rankings of the CommonsenseQA task when using the alphabet option format compared to the short text format (more experimental details can be found in Appendix D.1). This suggests that alphabet options may be unreliable, and we advocate for reducing the reliance on the alphabet option format in existing evaluation frameworks.
6 Limitations
Although xFinder has achieved encouraging results, we must acknowledge some limitations. Our smallest xFinder model, with only 500 million parameters, performs well as a key answer extraction component in evaluation tasks. Nevertheless, the deployment cost may still be high for some researchers. In the future, we promise to continuously update xFinder by exploring model compression techniques to support smaller versions, potentially enabling CPU deployment and thereby further enhancing xFinder’s flexibility.
7 Conclusion
In this paper, we analyzed the key factors in existing LLM evaluation frameworks that may lead to unreliable assessment results, highlighting the critical answer extraction module in LLM responses as a weak point. To address this issue, we introduce a key answer extraction model, xFinder, specifically for extracting key answers in LLM responses. Extensive experimental results demonstrate that xFinder, compared to existing RegEx-based key answer extraction methods, can extract key answers from LLM outputs more accurately and efficiently, thereby improving evaluation reliability. This study marks the first step in constructing a reliable evaluation framework. In the future, we will continue addressing other key issues in the reliable evaluation of LLMs, providing a foundation for accurately evaluating LLM performance.
Appendix A Datasets and Models
A.1 Datasets
Table 3 provides a summary of all the datasets utilized in our experiments. We undertook extensive preprocessing activities on the raw data from these datasets, ensuring that all samples were transformed into a consistent and unified format. From our original dataset, we sampled 2000 data points from the training set and 1000 data points from the test set (or dev set if a test set was unavailable). Subsequently, we categorized the questions into four types based on the key answer format in the dataset: alphabet, short text, categorical label, and math type.
Dataset Key Answer Type #Train #Test License
AgNews C 2000 1000 Unspecified
Amazon C 2000 1000 Apache-2.0
ARC-challenge A/T 2000 1000 CC BY-SA 4.0
ARC-easy A/T 1391 1000 CC BY-SA 4.0
BOOlQ C 2000 1000 CC BY-SA 3.0
ComonsenseQA A/T 2000 1000 MIT License
DBPedia C 2000 1000 CC-BY-SA-3.0
GSM8K M 2000 1000 MIT License
hellaswag A 2000 1000 MIT License
MATH M 1974 986 MIT License
MetaMathQA M 2000 1000 Apache-2.0
MMLU A/T 2000 638 MIT License
MultiArith M 420 180 Unspecified
OpenbookQA A/T 2000 500 Apache-2.0
QNLI C 2000 1000 Unspecified
SIQA A/T 2000 1000 Unspecified
Subj C 2000 1000 Unspecified
TREC C 1000 500 Unspecified
WiC C 2000 638 CC BY-NC 4.0
Table 3: Datasets Description. The datasets are sorted by their names in alphabetical order. Key Answer Type refers to the format of the key answer: A/T/C/M stands for alphabet / short text / categorical label / math type options respectively. Asterisk (*) denotes that the data within these datasets are utilized for fine-tuning our xFinder model.
A.2 Large Language Models
We selected the models listed in Table 4 from a vast array of open-source LLMs for our experiments. Among them, the specific version of GPT-4 we used for annotating the KAF dataset is GPT4-1106-preview. The training and test sets of the KAF dataset used to generate question-answer pairs include 10 LLMs: LLaMA2-7B-Chat, Baichuan2-7B-Chat, ChatGLM3-6B, Phi-2, InternLM2-7B-Chat, Qwen1.5-4B-Chat, Qwen1.5-1.8B-Chat, Qwen1.5-0.5B-Chat, Gemma-7B-it, and Gemma-2B-it. The generalization set used to generate question-answer pairs includes 8 LLMs: LLaMA3-8B-Instruct, LLaMA2-7B-Chat, ChatGLM3-6B, Phi-2, InternLM2-1.8B-Chat, Qwen1.5-14B-Chat, Qwen1.5-MoE-A2.7B-Chat, and Gemma-2B-it.
Model #Para. Type Publisher Date
LLaMA2 7B Chat Meta 2023.07
Baichuan2 7B Chat Baichuan Inc. 2023.09
ChatGLM3 6B Chat Tsinghua 2023.10
GPT4-1106-preview NaN Chat OpenAI 2023.11
Phi-2 2.7B Base Microsoft 2023.12
InternLM2 7B Chat ShLab 2024.01
InternLM2 1.8B Chat ShLab 2024.01
Qwen1.5 14B Chat Alibaba 2024.02
Qwen1.5 4B Chat Alibaba 2024.02
Qwen1.5 1.8B Chat Alibaba 2024.02
Qwen1.5 0.5B Chat Alibaba 2024.02
Gemma 7B Instruct Google 2024.02
Gemma 2B Instruct Google 2024.02
Qwen1.5-MoE 14.3B Chat Alibaba 2024.03
LLaMA3 8B Instruct Meta 2024.04
Table 4: LLMs sorted by release date. All LLMs are chat or instruct models. Asterisk (*) denotes estimated value, NaN denotes no public data available, and 175B denotes 175 billion.
A.3 Annotation Details
In our manual annotation project, all annotators are volunteers with a background in natural language processing, and each annotator holds at least a master’s degree. The project team consists of five annotators, with a gender ratio of three men to two women. Regarding their remuneration, they first receive the standard employee salary. We also pay them an additional 2 RMB per annotated data item, with the average time for annotating each task being about 20 seconds. Additionally, our engineering team participates in data review, primarily responsible for supervising and evaluating annotation quality. The entire annotation process lasted for 15 days.
Appendix B Fine-Tuning Hyperparameters
When fine-tuning xFinder using the XTuner framework and the QLoRA method, all base models were trained with identical hyperparameters. The training was conducted on 8x A100 GPUs. We focused on the hyperparameters specified in Table 5, while all other hyperparameters were set to their default values.
Hyperparameter Setting
Batch Size 1
Maximum Length 2048
Learning Rate 2e-4
Pack to Max Length True
Optimizer Type AdamW
Betas (0.9, 0.999)
Weight Decay 0
Warmup Ratio 0.03
Table 5: xFinder fine-tuning hyperparameter setting.
Appendix C Detailed Experimental Results
C.1 Extraction Accuracy Results
Table 6 presents the extraction accuracy of all 19 LLMs fine-tuned on the KAF generalization set. It can be observed that LLMs with different parameter magnitudes achieve stable and high accuracy levels after training on the KAF training set.
alphabet option short text categorical label math Overall
xFinder-qwen1505 0.9735 0.9683 0.9805 0.9276 0.9688
xFinder-qwen1505chat 0.9634 0.9635 0.9811 0.9351 0.9653
xFinder-qwen1518 0.9735 0.9781 0.9755 0.9457 0.9712
xFinder-qwen1518chat 0.9709 0.9708 0.9736 0.9382 0.9673
xFinder-internlm218 0.9608 0.9695 0.9667 0.9201 0.9587
xFinder-gemma2 0.9698 0.9744 0.9774 0.9276 0.9673
xFinder-gemma2it 0.9661 0.9756 0.9755 0.9246 0.9651
xFinder-qwen154 0.9730 0.9708 0.9698 0.9291 0.9657
xFinder-qwen154chat 0.9671 0.9793 0.9686 0.9306 0.9647
xFinder-chatglm36base 0.9735 0.9793 0.9730 0.9291 0.9684
xFinder-chatglm36 0.9698 0.9769 0.9742 0.9261 0.9665
xFinder-qwen157 0.9703 0.9744 0.9774 0.9306 0.9680
xFinder-qwen157chat 0.9735 0.9842 0.9748 0.9321 0.9702
xFinder-internlm27 0.9740 0.9756 0.9730 0.9336 0.9686
xFinder-baichuan27chat 0.9724 0.9756 0.9717 0.9276 0.9667
xFinder-gemma7 0.9751 0.9805 0.9736 0.9367 0.9704
xFinder-gemma7it 0.9682 0.9708 0.9226 0.9367 0.9498
xFinder-llama38 0.9698 0.9829 0.9692 0.9276 0.9661
xFinder-llama38it 0.9709 0.9683 0.9711 0.9382 0.9661
Table 6: Extraction accuracy of xFinder fine-tuned on all 19 LLMs on the KAF test set. The first column lists the models fine-tuned using the KAF training set, including Qwen1.5-0.5B, Qwen1.5-0.5B-Chat, Qwen1.5-1.8B, Qwen1.5-1.8B-Chat, InternLM2-1.8B, Gemma-2B, Gemma-2B-it, Qwen1.5-4B, Qwen1.5-4B-Chat, ChatGLM3-6B-base, ChatGLM3-6B, Qwen1.5-7B, Qwen1.5-7B-Chat, InternLM2-7B, Baichuan2-7B-Chat, Gemma-7B, Gemma-7B-it, Llama-3-8B, and Llama-3-8B-Instruct models.
Table 7 displays the extraction accuracy, $\Delta_{acc}$, and $\Delta_{acc}/N$ of all 19 LLMs fine-tuned on the KAF generalization set.
alphabet option short text categorical label math Overall $\Delta_{acc}$ $\Delta_{acc}/N$
xFinder-qwen1505 0.9477 0.9335 0.9281 0.9234 0.9342 0.0277 0.0554
xFinder-qwen1505chat 0.9289 0.9283 0.9163 0.9375 0.9255 0.0190 0.0380
xFinder-qwen1518 0.9508 0.9439 0.9406 0.9500 0.9456 0.0391 0.0217
xFinder-qwen1518chat 0.9508 0.9356 0.9181 0.9531 0.9362 0.0297 0.0165
xFinder-internlm218 0.9500 0.9407 0.8400 0.9344 0.9065 0 0
xFinder-gemma2 0.9484 0.9335 0.9337 0.9437 0.9393 0.0328 0.0164
xFinder-gemma2it 0.9344 0.9262 0.9263 0.9391 0.9304 0.0239 0.0120
xFinder-qwen154 0.9484 0.9345 0.9387 0.9484 0.9420 0.0355 0.0088
xFinder-qwen154chat 0.9492 0.9407 0.9237 0.9531 0.9389 0.0324 0.0081
xFinder-chatglm36base 0.9484 0.9324 0.9481 0.9484 0.9449 0.0384 0.0064
xFinder-chatglm36 0.9563 0.9376 0.9163 0.9484 0.9369 0.0304 0.0051
xFinder-qwen157 0.9570 0.9345 0.9519 0.9391 0.9478 0.0413 0.0059
xFinder-qwen157chat 0.9563 0.9345 0.9319 0.9422 0.9409 0.0344 0.0049
xFinder-internlm27 0.9563 0.9563 0.9281 0.9375 0.9436 0.0371 0.0053
xFinder-baichuan27chat 0.9477 0.9345 0.9431 0.9313 0.9409 0.0344 0.0049
xFinder-gemma7 0.9477 0.9345 0.9456 0.9500 0.9444 0.0379 0.0054
xFinder-gemma7it 0.9398 0.9200 0.8938 0.9469 0.9201 0.0136 0.0019
xFinder-llama38 0.9563 0.9376 0.9469 0.9516 0.9482 0.0417 0.0052
xFinder-llama38it 0.9547 0.9428 0.9537 0.9547 0.9518 0.0453 0.0057
Table 7: Extraction accuracy of xFinder fine-tuned on all 19 LLMs on the KAF generalization set, along with $\Delta_{acc}$ and $\Delta_{acc}/N$ values. The first column lists the models fine-tuned using the KAF training set: Qwen1.5-0.5B, Qwen1.5-0.5B-Chat, Qwen1.5-1.8B, Qwen1.5-1.8B-Chat, InternLM2-1.8B, Gemma-2B, Gemma-2B-it, Qwen1.5-4B, Qwen1.5-4B-Chat, ChatGLM3-6B-base, ChatGLM3-6B, Qwen1.5-7B, Qwen1.5-7B-Chat, InternLM2-7B, Baichuan2-7B-Chat, Gemma-7B, Gemma-7B-it, Llama-3-8B, and Llama-3-8B-Instruct.
C.2 Evaluation in Real-World Scenarios Results
C.2.1 Alphabet & Short Text Option Tasks
LLMs OpenCompass UltraEval LM Eval Harness xFinder-llama38it xFinder-qwen1505
Baichuan2-7B-Chat 39.17% 38.15% 52.80% 54.93% 55.04%
ChatGLM3-6B 56.56% 44.46% 56.26% 57.27% 57.17%
Gemm-2B-it 43.23% 42.22% 43.23% 44.56% 43.44%
Gemma-7B-it 50.76% 45.88% 35.10% 71.01% 70.80%
InternLM2-1.8B-Chat 59.31% 59.21% 58.70% 59.51% 59.61%
InternLM2-7B-Chat 76.70% 77.72% 77.72% 77.92% 77.72%
Llama3-8B-Instruct 78.03% 77.92% 77.92% 78.03% 78.03%
Phi-2 66.12% 65.82% 70.91% 70.91% 71.01%
Qwen1.5-14B-Chat 83.32% 64.29% 76.60% 85.96% 86.06%
Qwen1.5-MoE-A2.7B-Chat 70.30% 67.96% 70.19% 73.35% 73.45%
Table 8: Evaluation results of LLMs on the ARC-c (alphabet) task, presenting accuracy scores using different key answer extraction methods.
LLMs OpenCompass UltraEval LM Eval Harness xFinder-llama38it xFinder-qwen1505
Baichuan2-7B-Chat 40.80% 40.80% 47.50% 48.40% 48.40%
ChatGLM3-6B 26.70% 26.70% 37.30% 40.70% 39.70%
Gemm-2B-it 8.00% 8.00% 25.10% 31.70% 29.60%
Gemma-7B-it 31.80% 31.80% 58.10% 65.20% 63.90%
InternLM2-1.8B-Chat 34.10% 34.10% 43.80% 44.30% 42.20%
InternLM2-7B-Chat 57.20% 57.20% 70.40% 71.00% 70.00%
Llama3-8B-Instruct 60.50% 60.50% 75.10% 74.80% 74.70%
Phi-2 41.30% 41.30% 48.40% 49.80% 49.50%
Qwen1.5-14B-Chat 60.10% 60.10% 77.80% 80.60% 80.60%
Qwen1.5-MoE-A2.7B-Chat 41.30% 41.30% 61.10% 63.20% 62.50%
Table 9: Evaluation results of LLMs on the ARC-c (text) task, presenting accuracy scores using different key answer extraction methods.
LLMs OpenCompass UltraEval LM Eval Harness xFinder-llama38it xFinder-qwen1505
Baichuan2-7B-Chat 44.30% 42.70% 71.20% 74.10% 73.60%
ChatGLM3-6B 71.40% 47.50% 70.60% 71.90% 72.10%
Gemm-2B-it 61.70% 59.50% 62.00% 62.80% 61.90%
Gemma-7B-it 58.80% 54.80% 39.20% 84.40% 84.10%
InternLM2-1.8B-Chat 74.90% 75.60% 76.00% 76.40% 76.10%
InternLM2-7B-Chat 89.40% 90.50% 90.20% 90.40% 90.50%
Llama3-8B-Instruct 92.70% 92.70% 92.70% 92.70% 92.70%
Phi-2 78.00% 77.70% 84.20% 84.60% 84.60%
Qwen1.5-14B-Chat 93.40% 76.70% 88.20% 95.60% 95.60%
Qwen1.5-MoE-A2.7B-Chat 83.40% 81.20% 84.20% 86.10% 85.80%
Table 10: Evaluation results of LLMs on the ARC-e (alphabet) task, presenting accuracy scores using different key answer extraction methods.
LLMs OpenCompass UltraEval LM Eval Harness xFinder-llama38it xFinder-qwen1505
Baichuan2-7B-Chat 61.40% 61.40% 65.20% 67.00% 67.20%
ChatGLM3-6B 39.60% 39.60% 49.20% 52.70% 51.60%
Gemm-2B-it 8.20% 8.20% 32.80% 45.80% 41.20%
Gemma-7B-it 51.80% 51.80% 76.80% 82.70% 81.60%
InternLM2-1.8B-Chat 53.30% 53.30% 60.70% 60.50% 58.70%
InternLM2-7B-Chat 79.40% 79.40% 86.50% 86.90% 86.20%
Llama3-8B-Instruct 80.40% 80.40% 89.90% 89.70% 89.40%
Phi-2 62.80% 62.80% 67.30% 69.50% 69.30%
Qwen1.5-14B-Chat 74.10% 74.10% 88.50% 91.90% 91.80%
Qwen1.5-MoE-A2.7B-Chat 56.70% 56.70% 75.80% 78.60% 77.60%
Table 11: Evaluation results of LLMs on the ARC-e (text) task, presenting accuracy scores using different key answer extraction methods.
LLMs OpenCompass UltraEval LM Eval Harness xFinder-llama38it xFinder-qwen1505
Baichuan2-7B-Chat 36.00% 27.50% 58.60% 60.00% 59.90%
ChatGLM3-6B 61.20% 31.80% 57.00% 62.30% 61.70%
Gemm-2B-it 44.70% 39.90% 44.90% 45.90% 45.00%
Gemma-7B-it 55.10% 34.90% 41.00% 69.90% 69.90%
InternLM2-1.8B-Chat 71.50% 71.40% 73.70% 73.10% 72.90%
InternLM2-7B-Chat 82.60% 66.90% 83.70% 83.70% 83.80%
Llama3-8B-Instruct 71.60% 58.30% 71.70% 71.30% 71.30%
Phi-2 59.50% 51.90% 61.30% 61.40% 61.30%
Qwen1.5-14B-Chat 83.40% 60.20% 74.80% 83.60% 83.40%
Qwen1.5-MoE-A2.7B-Chat 73.00% 57.80% 75.70% 77.70% 77.50%
Table 12: Evaluation results of LLMs on the CommonsenseQA (alphabet) task, presenting accuracy scores using different key answer extraction methods.
LLMs OpenCompass UltraEval LM Eval Harness xFinder-llama38it xFinder-qwen1505
Baichuan2-7B-Chat 51.20% 51.20% 51.50% 52.40% 52.20%
ChatGLM3-6B 31.80% 31.80% 41.90% 46.60% 46.60%
Gemm-2B-it 5.70% 5.70% 31.40% 39.10% 38.10%
Gemma-7B-it 53.20% 53.20% 63.80% 68.00% 67.60%
InternLM2-1.8B-Chat 58.00% 58.00% 58.60% 59.00% 58.60%
InternLM2-7B-Chat 74.80% 74.80% 75.80% 78.50% 78.00%
Llama3-8B-Instruct 70.20% 70.20% 71.40% 71.20% 71.50%
Phi-2 41.80% 41.80% 42.40% 42.60% 42.60%
Qwen1.5-14B-Chat 70.50% 70.50% 75.60% 78.60% 78.90%
Qwen1.5-MoE-A2.7B-Chat 55.50% 55.50% 67.80% 70.40% 70.20%
Table 13: Evaluation results of LLMs on the CommonsenseQA (text) task, presenting accuracy scores using different key answer extraction methods.
LLMs OpenCompass UltraEval LM Eval Harness xFinder-llama38it xFinder-qwen1505
Baichuan2-7B-Chat 36.60% 34.80% 41.30% 44.40% 44.30%
ChatGLM3-6B 39.10% 31.20% 39.20% 39.40% 39.60%
Gemm-2B-it 37.40% 36.50% 37.50% 38.10% 37.30%
Gemma-7B-it 40.80% 35.40% 24.90% 50.00% 49.70%
InternLM2-1.8B-Chat 42.10% 44.30% 45.00% 45.80% 45.10%
InternLM2-7B-Chat 54.10% 54.80% 55.00% 55.70% 55.00%
Llama3-8B-Instruct 62.50% 62.00% 62.70% 62.60% 62.50%
Phi-2 46.10% 45.20% 48.30% 48.80% 48.60%
Qwen1.5-14B-Chat 65.40% 49.70% 59.80% 66.00% 66.00%
Qwen1.5-MoE-A2.7B-Chat 50.80% 48.60% 51.60% 54.50% 54.20%
Table 14: Evaluation results of LLMs on the MMLU (alphabet) task, presenting accuracy scores using different key answer extraction methods.
LLMs OpenCompass UltraEval LM Eval Harness xFinder-llama38it xFinder-qwen1505
Baichuan2-7B-Chat 26.00% 26.00% 37.20% 35.70% 35.90%
ChatGLM3-6B 13.40% 13.40% 24.00% 22.70% 22.40%
Gemm-2B-it 5.20% 5.20% 23.20% 24.50% 21.50%
Gemma-7B-it 19.60% 19.60% 44.00% 42.30% 41.70%
InternLM2-1.8B-Chat 22.30% 22.30% 33.30% 29.50% 29.80%
InternLM2-7B-Chat 30.10% 30.10% 48.90% 44.60% 44.80%
Llama3-8B-Instruct 36.80% 36.80% 55.20% 49.60% 49.80%
Phi-2 22.30% 22.30% 32.90% 30.60% 30.70%
Qwen1.5-14B-Chat 31.60% 31.60% 53.10% 52.60% 51.90%
Qwen1.5-MoE-A2.7B-Chat 25.60% 25.60% 43.60% 41.80% 41.30%
Table 15: Evaluation results of LLMs on the MMLU (text) task, presenting accuracy scores using different key answer extraction methods.
LLMs OpenCompass UltraEval LM Eval Harness xFinder-llama38it xFinder-qwen1505
Baichuan2-7B-Chat 39.00% 37.00% 47.80% 50.40% 50.40%
ChatGLM3-6B 52.80% 41.80% 52.00% 53.20% 53.20%
Gemm-2B-it 51.80% 51.00% 51.80% 52.60% 51.80%
Gemma-7B-it 57.40% 28.60% 49.40% 65.20% 65.00%
InternLM2-1.8B-Chat 55.40% 56.80% 57.20% 56.20% 57.20%
InternLM2-7B-Chat 77.60% 78.20% 78.00% 78.40% 78.60%
Llama3-8B-Instruct 74.60% 74.60% 74.40% 74.60% 74.60%
Phi-2 65.40% 64.60% 68.00% 68.00% 68.00%
Qwen1.5-14B-Chat 83.00% 68.20% 75.40% 84.20% 84.20%
Qwen1.5-MoE-A2.7B-Chat 69.80% 66.60% 69.20% 72.00% 71.60%
Table 16: Evaluation results of LLMs on the OpenbookQA (alphabet) task, presenting accuracy scores using different key answer extraction methods.
LLMs OpenCompass UltraEval LM Eval Harness xFinder-llama38it xFinder-qwen1505
Baichuan2-7B-Chat 41.60% 41.60% 44.20% 45.40% 45.60%
ChatGLM3-6B 22.60% 22.60% 34.20% 36.20% 36.20%
Gemm-2B-it 5.20% 5.20% 24.00% 30.80% 30.60%
Gemma-7B-it 42.60% 42.60% 58.60% 63.40% 63.00%
InternLM2-1.8B-Chat 38.40% 38.40% 44.00% 43.80% 42.20%
InternLM2-7B-Chat 61.40% 61.40% 70.20% 70.00% 70.00%
Llama3-8B-Instruct 67.80% 67.80% 72.80% 72.80% 72.60%
Phi-2 34.60% 34.60% 36.60% 37.40% 37.40%
Qwen1.5-14B-Chat 68.00% 68.00% 75.40% 81.00% 80.80%
Qwen1.5-MoE-A2.7B-Chat 39.60% 39.60% 54.20% 55.40% 55.20%
Table 17: Evaluation results of LLMs on the OpenbookQA (text) task, presenting accuracy scores using different key answer extraction methods.
LLMs OpenCompass UltraEval LM Eval Harness xFinder-llama38it xFinder-qwen1505
Baichuan2-7B-Chat 44.50% 43.50% 58.90% 60.40% 60.20%
ChatGLM3-6B 66.10% 60.30% 66.00% 66.50% 66.50%
Gemm-2B-it 53.00% 52.70% 53.00% 53.40% 53.00%
Gemma-7B-it 63.00% 15.10% 55.40% 66.00% 66.30%
InternLM2-1.8B-Chat 69.20% 69.80% 70.80% 69.80% 70.00%
InternLM2-7B-Chat 75.90% 77.30% 77.70% 77.10% 77.20%
Llama3-8B-Instruct 71.00% 70.50% 70.90% 70.60% 70.70%
Phi-2 63.60% 63.60% 64.00% 64.00% 64.00%
Qwen1.5-14B-Chat 79.00% 58.30% 62.40% 78.90% 79.10%
Qwen1.5-MoE-A2.7B-Chat 68.90% 67.30% 71.60% 72.50% 72.50%
Table 18: Evaluation results of LLMs on the SIQA (alphabet) task, presenting accuracy scores using different key answer extraction methods.
LLMs OpenCompass UltraEval LM Eval Harness xFinder-llama38it xFinder-qwen1505
Baichuan2-7B-Chat 50.00% 50.00% 52.90% 53.70% 53.50%
ChatGLM3-6B 32.40% 32.40% 49.70% 52.50% 52.90%
Gemm-2B-it 8.40% 8.40% 35.20% 45.30% 43.90%
Gemma-7B-it 32.50% 32.50% 62.60% 64.00% 64.30%
InternLM2-1.8B-Chat 53.10% 53.10% 60.20% 59.60% 59.80%
InternLM2-7B-Chat 58.40% 58.40% 68.60% 67.70% 68.00%
Llama3-8B-Instruct 61.30% 61.30% 64.90% 64.60% 64.70%
Phi-2 43.50% 43.50% 44.20% 45.80% 45.70%
Qwen1.5-14B-Chat 62.50% 62.50% 68.80% 72.90% 72.70%
Qwen1.5-MoE-A2.7B-Chat 48.40% 48.40% 62.70% 66.10% 65.40%
Table 19: Evaluation results of LLMs on the SIQA (text) task, presenting accuracy scores using different key answer extraction methods.
LLMs OpenCompass UltraEval LM Eval Harness xFinder-llama38it xFinder-qwen1505
Baichuan2-7B-Chat 46.40% 45.80% 46.80% 47.20% 46.70%
ChatGLM3-6B 44.50% 43.60% 43.90% 44.20% 43.80%
Gemm-2B-it 29.80% 27.80% 29.80% 29.80% 29.70%
Gemma-7B-it 50.20% 21.30% 42.30% 51.20% 51.10%
InternLM2-1.8B-Chat 61.20% 62.20% 60.30% 61.80% 61.30%
InternLM2-7B-Chat 73.40% 72.70% 73.40% 73.40% 73.20%
Llama3-8B-Instruct 69.30% 68.80% 69.30% 68.80% 68.80%
Phi-2 34.40% 34.20% 34.40% 34.30% 34.20%
Qwen1.5-14B-Chat 78.50% 75.60% 75.50% 78.60% 78.60%
Qwen1.5-MoE-A2.7B-Chat 62.60% 59.70% 63.60% 66.10% 65.50%
Table 20: Evaluation results of LLMs on the hellaswag (alphabet) task, presenting accuracy scores using different key answer extraction methods.
C.2.2 Categorical Label Option Tasks
LLMs OpenCompass UltraEval LM Eval Harness xFinder-llama38it xFinder-qwen1505
Baichuan2-7B-Chat 83.40% 83.40% 83.60% 83.70% 83.60%
ChatGLM3-6B 86.50% 86.50% 95.20% 95.20% 95.20%
Gemm-2B-it 81.50% 81.50% 85.30% 85.10% 85.30%
Gemma-7B-it 73.60% 73.60% 94.10% 94.30% 94.20%
InternLM2-1.8B-Chat 96.40% 96.40% 96.60% 96.50% 95.50%
InternLM2-7B-Chat 96.10% 96.10% 96.20% 96.10% 96.10%
Llama3-8B-Instruct 93.40% 93.40% 93.40% 93.40% 93.40%
Phi-2 91.60% 91.60% 91.80% 91.80% 91.80%
Qwen1.5-14B-Chat 95.50% 95.50% 95.80% 95.50% 95.70%
Qwen1.5-MoE-A2.7B-Chat 90.50% 90.50% 93.70% 92.50% 92.60%
Table 21: Evaluation results of LLMs on the Amazon task, presenting accuracy scores using different key answer extraction methods.
LLMs OpenCompass UltraEval LM Eval Harness xFinder-llama38it xFinder-qwen1505
Baichuan2-7B-Chat 63.60% 63.60% 63.60% 63.60% 63.60%
ChatGLM3-6B 20.60% 20.60% 60.50% 60.20% 61.00%
Gemm-2B-it 24.80% 24.80% 63.50% 63.20% 63.80%
Gemma-7B-it 68.60% 68.60% 69.40% 69.70% 69.70%
InternLM2-1.8B-Chat 55.00% 55.00% 64.60% 61.50% 58.20%
InternLM2-7B-Chat 76.70% 76.70% 76.70% 76.50% 76.50%
Llama3-8B-Instruct 68.60% 68.60% 68.50% 68.60% 68.60%
Phi-2 64.10% 64.10% 64.10% 64.10% 64.10%
Qwen1.5-14B-Chat 41.90% 41.90% 74.60% 74.90% 74.90%
Qwen1.5-MoE-A2.7B-Chat 24.30% 24.30% 63.60% 63.50% 63.60%
Table 22: Evaluation results of LLMs on the BoolQ task, presenting accuracy scores using different key answer extraction methods.
LLMs OpenCompass UltraEval LM Eval Harness xFinder-llama38it xFinder-qwen1505
Baichuan2-7B-Chat 58.40% 58.40% 58.40% 58.50% 58.50%
ChatGLM3-6B 42.60% 42.60% 63.50% 64.90% 65.50%
Gemm-2B-it 54.30% 54.30% 55.40% 55.40% 55.20%
Gemma-7B-it 53.30% 53.30% 56.80% 56.80% 57.70%
InternLM2-1.8B-Chat 51.50% 51.50% 69.50% 67.60% 68.40%
InternLM2-7B-Chat 60.80% 60.80% 61.60% 60.80% 67.80%
Llama3-8B-Instruct 72.10% 72.10% 72.10% 72.10% 72.20%
Phi-2 48.40% 48.40% 48.40% 48.40% 48.40%
Qwen1.5-14B-Chat 37.20% 37.20% 79.70% 78.00% 79.60%
Qwen1.5-MoE-A2.7B-Chat 35.80% 35.80% 77.00% 76.80% 77.10%
Table 23: Evaluation results of LLMs on the QNLI task, presenting accuracy scores using different key answer extraction methods.
LLMs OpenCompass UltraEval LM Eval Harness xFinder-llama38it xFinder-qwen1505
Baichuan2-7B-Chat 53.92% 53.92% 53.92% 53.92% 53.92%
ChatGLM3-6B 32.29% 32.29% 51.88% 52.35% 52.35%
Gemm-2B-it 54.39% 54.39% 54.39% 54.39% 54.39%
Gemma-7B-it 50.31% 50.31% 50.31% 50.47% 50.47%
InternLM2-1.8B-Chat 54.08% 54.08% 54.08% 53.76% 54.08%
InternLM2-7B-Chat 51.57% 51.57% 51.57% 51.57% 51.57%
Llama3-8B-Instruct 56.90% 56.90% 56.90% 56.90% 56.90%
Phi-2 50.31% 50.31% 50.31% 50.16% 50.31%
Qwen1.5-14B-Chat 47.65% 47.65% 64.42% 64.42% 64.42%
Qwen1.5-MoE-A2.7B-Chat 45.45% 45.45% 57.99% 57.84% 57.84%
Table 24: Evaluation results of LLMs on the WiC task, presenting accuracy scores using different key answer extraction methods.
C.2.3 Math Option Tasks
LLMs OpenCompass UltraEval LM Eval Harness xFinder-llama38it xFinder-qwen1505
Baichuan2-7B-Chat 6.60% 5.50% 5.50% 6.70% 6.80%
ChatGLM3-6B 43.60% 0.00% 37.10% 43.60% 43.60%
Gemm-2B-it 7.90% 3.10% 3.20% 8.70% 7.90%
Gemma-7B-it 35.70% 1.20% 33.30% 36.80% 36.60%
InternLM2-1.8B-Chat 14.00% 2.40% 23.90% 31.80% 26.20%
InternLM2-7B-Chat 49.50% 5.10% 50.80% 63.10% 58.60%
Llama3-8B-Instruct 77.00% 2.20% 80.30% 80.30% 80.20%
Phi-2 15.60% 12.40% 13.40% 16.80% 16.50%
Qwen1.5-14B-Chat 74.40% 0.40% 76.00% 77.00% 76.70%
Qwen1.5-MoE-A2.7B-Chat 46.80% 3.80% 42.10% 48.00% 48.20%
Table 25: Evaluation results of LLMs on the GSM8K task, presenting accuracy scores using different key answer extraction methods.
LLMs OpenCompass UltraEval LM Eval Harness xFinder-llama38it xFinder-qwen1505
Baichuan2-7B-Chat 6.39% 6.39% 5.68% 7.20% 7.10%
ChatGLM3-6B 23.02% 8.22% 4.56% 23.02% 22.82%
Gemm-2B-it 6.39% 4.56% 5.07% 8.82% 8.82%
Gemma-7B-it 13.18% 1.72% 11.46% 14.40% 14.10%
InternLM2-1.8B-Chat 11.16% 3.55% 8.32% 11.66% 10.55%
InternLM2-7B-Chat 19.57% 4.67% 16.23% 21.20% 20.39%
Llama3-8B-Instruct 30.73% 7.10% 24.44% 30.63% 30.73%
Phi-2 9.03% 8.82% 6.90% 10.24% 9.33%
Qwen1.5-14B-Chat 31.34% 4.26% 26.06% 36.71% 36.31%
Qwen1.5-MoE-A2.7B-Chat 17.65% 7.91% 6.69% 22.21% 22.52%
Table 26: Evaluation results of LLMs on the MATH task, presenting accuracy scores using different key answer extraction methods.
LLMs OpenCompass UltraEval LM Eval Harness xFinder-llama38it xFinder-qwen1505
Baichuan2-7B-Chat 5.00% 3.33% 3.33% 5.00% 5.00%
ChatGLM3-6B 72.78% 1.11% 63.33% 72.78% 72.78%
Gemm-2B-it 23.89% 4.44% 4.44% 26.11% 23.89%
Gemma-7B-it 84.44% 1.67% 77.78% 84.44% 83.33%
InternLM2-1.8B-Chat 30.56% 16.67% 36.67% 59.44% 45.56%
InternLM2-7B-Chat 56.67% 36.67% 67.78% 81.67% 70.00%
Llama3-8B-Instruct 97.78% 1.67% 97.78% 97.78% 97.78%
Phi-2 33.33% 29.44% 28.33% 41.67% 38.89%
Qwen1.5-14B-Chat 98.33% 0.56% 98.33% 98.33% 98.33%
Qwen1.5-MoE-A2.7B-Chat 62.78% 7.78% 52.78% 63.33% 63.33%
Table 27: Evaluation results of LLMs on the MultiArith task, presenting accuracy scores using different key answer extraction methods.
Appendix D Bump Charts of All Tasks
D.1 Alphabet & Short Text Option Tasks
Figure 5: Bump Chart of ARC-c
Figure 6: Bump Chart of ARC-e
Figure 7: Bump Chart of CommonsenseQA
Figure 8: Bump Chart of MMLU
Figure 9: Bump Chart of OpenbookQA
Figure 10: Bump Chart of SIQA
Figure 11: Bump Chart of hellaswag (alphabet)
D.2 Categorical Label Option Tasks
Figure 12: Bump Chart of Amazon
Figure 13: Bump Chart of BoolQ
Figure 14: Bump Chart of QNLI
Figure 15: Bump Chart of WiC
D.3 Math Option Tasks
Figure 16: Bump Chart of GSM8K
Figure 17: Bump Chart of MATH
Figure 18: Bump Chart of MultiArith
Appendix E Prompts
E.1 Prompts for LLM Responses Generation & Real-World Scenarios Evaluation
In these figures, the red part represents the System prompt, the blue part represents the Demonstration, and the black part represents the Question.
You are an expert in commonsense reasoning, Please choose the answer from options A to D, corresponding to the question. Q: Coal is a fossil fuel that is formed from Answer Choices: (A) water eroding the land. (B) meteors hitting Earth. (C) repeated volcanic eruptions. (D) the decay of organic material. A:
Figure 19: 0-shot prompt for LLM response generation (using alpha option type question as an example).
You are an expert in commonsense reasoning, Please choose the answer from options A to D, corresponding to the question. Q: Coal is a fossil fuel that is formed from Answer Choices: (A) water eroding the land. (B) meteors hitting Earth. (C) repeated volcanic eruptions. (D) the decay of organic material. A: End your final answer with ’The answer is <answer>.’
Figure 20: 0-shot-restrict prompt for LLM response generation & Real-World scenarios evaluation (using alpha option type question as an example).
You are an expert in commonsense reasoning, Please choose the answer from options A to D, corresponding to the question. Q: Coal is a fossil fuel that is formed from Answer Choices: (A) water eroding the land. (B) meteors hitting Earth. (C) repeated volcanic eruptions. (D) the decay of organic material. A: Let’s think step by step.
Figure 21: 0-shot-cot prompt for LLM response generation (using alpha option type question as an example).
You are an expert in commonsense reasoning, Please choose the answer from options A to D, corresponding to the question. Q: Coal is a fossil fuel that is formed from Answer Choices: (A) water eroding the land. (B) meteors hitting Earth. (C) repeated volcanic eruptions. (D) the decay of organic material. A: End your final answer with ’The answer is <answer>.’ Let’s think step by step.
Figure 22: 0-shot-cot-restrict prompt for LLM response generation (using alpha option type question as an example).
You are an expert in commonsense reasoning, Please choose the answer from options A to D, corresponding to the question. ***** Start In-Context Examples ***** Q: Which of the following helps to produce urine in humans and other mammals? Answer Choices: (A) bladder (B) urethra (C) kidneys (D) ureter A: The answer is (C). Q: An atom consists of a nucleus surrounded by Answer Choices: (A) ions. (B) protons. (C) neutrons. (D) electrons. A: The answer is (D). Q: Which lists the diameter of the planets in order from smallest to largest? Answer Choices: (A) Venus, Earth, Mercury, Mars (B) Earth, Mars, Venus, Mercury (C) Mars, Mercury, Earth, Venus (D) Mercury, Mars, Venus, Earth A: The answer is (D). Q: In a container, a mixture of water and salt is stirred so that the salt dissolves completely. Sand is added to this solution and allowed to settle to the bottom of the container. If the container is placed on a heat source and the liquid evaporates completely, what will be left in the container? Answer Choices: (A) Nothing will remain in the container. (B) Only salt will remain in the container. (C) Only sand will remain in the container. (D) Salt and sand will both remain in the container. A: The answer is (D). Q: Which instrument measures atmospheric pressure? Answer Choices: (A) barometer (B) hygrometer (C) thermometer (D) magnetometer A: The answer is (A). ***** End In-Context Examples ***** Q: Coal is a fossil fuel that is formed from Answer Choices: (A) water eroding the land. (B) meteors hitting Earth. (C) repeated volcanic eruptions. (D) the decay of organic material. A:
Figure 23: 5-shot prompt for LLM response generation (using alpha option type question as an example).
You are an expert in commonsense reasoning, Please choose the answer from options A to D, corresponding to the question. ***** Start In-Context Examples ***** Q: Which of the following helps to produce urine in humans and other mammals? Answer Choices: (A) bladder (B) urethra (C) kidneys (D) ureter A: The answer is (C). Q: An atom consists of a nucleus surrounded by Answer Choices: (A) ions. (B) protons. (C) neutrons. (D) electrons. A: The answer is (D). Q: Which lists the diameter of the planets in order from smallest to largest? Answer Choices: (A) Venus, Earth, Mercury, Mars (B) Earth, Mars, Venus, Mercury (C) Mars, Mercury, Earth, Venus (D) Mercury, Mars, Venus, Earth A: The answer is (D). Q: In a container, a mixture of water and salt is stirred so that the salt dissolves completely. Sand is added to this solution and allowed to settle to the bottom of the container. If the container is placed on a heat source and the liquid evaporates completely, what will be left in the container? Answer Choices: (A) Nothing will remain in the container. (B) Only salt will remain in the container. (C) Only sand will remain in the container. (D) Salt and sand will both remain in the container. A: The answer is (D). Q: Which instrument measures atmospheric pressure? Answer Choices: (A) barometer (B) hygrometer (C) thermometer (D) magnetometer A: The answer is (A). ***** End In-Context Examples ***** Q: Coal is a fossil fuel that is formed from Answer Choices: (A) water eroding the land. (B) meteors hitting Earth. (C) repeated volcanic eruptions. (D) the decay of organic material. A: End your final answer with ’The answer is <answer>.’
Figure 24: 5-shot-restrict prompt for LLM response generation (using alpha option type question as an example).
You are an expert in commonsense reasoning, Please choose the answer from options A to D, corresponding to the question. ***** Start In-Context Examples ***** Q: Which of the following helps to produce urine in humans and other mammals? Answer Choices: (A) bladder (B) urethra (C) kidneys (D) ureter A: The answer is (C). Q: An atom consists of a nucleus surrounded by Answer Choices: (A) ions. (B) protons. (C) neutrons. (D) electrons. A: The answer is (D). Q: Which lists the diameter of the planets in order from smallest to largest? Answer Choices: (A) Venus, Earth, Mercury, Mars (B) Earth, Mars, Venus, Mercury (C) Mars, Mercury, Earth, Venus (D) Mercury, Mars, Venus, Earth A: The answer is (D). Q: In a container, a mixture of water and salt is stirred so that the salt dissolves completely. Sand is added to this solution and allowed to settle to the bottom of the container. If the container is placed on a heat source and the liquid evaporates completely, what will be left in the container? Answer Choices: (A) Nothing will remain in the container. (B) Only salt will remain in the container. (C) Only sand will remain in the container. (D) Salt and sand will both remain in the container. A: The answer is (D). Q: Which instrument measures atmospheric pressure? Answer Choices: (A) barometer (B) hygrometer (C) thermometer (D) magnetometer A: The answer is (A). ***** End In-Context Examples ***** Q: Coal is a fossil fuel that is formed from Answer Choices: (A) water eroding the land. (B) meteors hitting Earth. (C) repeated volcanic eruptions. (D) the decay of organic material. A: Let’s think step by step.
Figure 25: 5-shot-cot prompt for LLM response generation (using alpha option type question as an example).
You are an expert in commonsense reasoning, Please choose the answer from options A to D, corresponding to the question. ***** Start In-Context Examples ***** Q: Which of the following helps to produce urine in humans and other mammals? Answer Choices: (A) bladder (B) urethra (C) kidneys (D) ureter A: The answer is (C). Q: An atom consists of a nucleus surrounded by Answer Choices: (A) ions. (B) protons. (C) neutrons. (D) electrons. A: The answer is (D). Q: Which lists the diameter of the planets in order from smallest to largest? Answer Choices: (A) Venus, Earth, Mercury, Mars (B) Earth, Mars, Venus, Mercury (C) Mars, Mercury, Earth, Venus (D) Mercury, Mars, Venus, Earth A: The answer is (D). Q: In a container, a mixture of water and salt is stirred so that the salt dissolves completely. Sand is added to this solution and allowed to settle to the bottom of the container. If the container is placed on a heat source and the liquid evaporates completely, what will be left in the container? Answer Choices: (A) Nothing will remain in the container. (B) Only salt will remain in the container. (C) Only sand will remain in the container. (D) Salt and sand will both remain in the container. A: The answer is (D). Q: Which instrument measures atmospheric pressure? Answer Choices: (A) barometer (B) hygrometer (C) thermometer (D) magnetometer A: The answer is (A). ***** End In-Context Examples ***** Q: Coal is a fossil fuel that is formed from Answer Choices: (A) water eroding the land. (B) meteors hitting Earth. (C) repeated volcanic eruptions. (D) the decay of organic material. A: End your final answer with ’The answer is <answer>.’ Let’s think step by step.
Figure 26: 5-shot-cot-restrict prompt for LLM response generation (using alpha option type question as an example).
E.2 Prompts for Auto Labeling Using GPT-4
As shown in Step 3 of Figure 3, we employed a self-consistency strategy for the automatic annotation using GPT-4. Specifically, we used two forms of prompts: a normal form and an XML form, as shown in Figure 27 and Figure 28, respectively. In these figures, the red part represents the System prompt, the orange part represents the Instruction prompt, and the black part represents the Question.
You are a help assistant tasked with extracting the precise key answer from given output sentences. You must only provide the extracted key answer without including any additional text. I will provide you with a question, output sentences along with an answer range. The output sentences are the response of the question provided. The answer range could either describe the type of answer expected or list all possible valid answers. Using the information provided, you must accurately and precisely determine and extract the intended key answer from the output sentences. Please don’t have your subjective thoughts about the question. If the output sentences present multiple different answers, carefully determine if the later provided answer is a correction or modification of a previous one. If so, extract this corrected or modified answer as the final response. Conversely, if the output sentences fluctuate between multiple answers without a clear final answer, you should output [No valid answer]. For a question of type "classify_text", if the output sentences do not clearly provide any answer within the specified answer range, then the output should be [No valid answer]. Additionally, if the answer range is a list and the answer given in the output sentences is not included in the answer range, also output [No valid answer]. Question: {question} Output sentences: {LLM response} Key answer type: {key answer type} Answer range: {answer range} Key extracted answer:
Figure 27: The prompt for automatic annotation by GPT-4: normal form.
You are a help assistant tasked with extracting the precise key answer from given output sentences. You must only provide the extracted key answer without including any additional text. I will provide you with a question, output sentences along with an answer range. The output sentences are the response of the question provided. The answer range could either describe the type of answer expected or list all possible valid answers. Using the information provided, you must accurately and precisely determine and extract the intended key answer from the output sentences. Please don’t have your subjective thoughts about the question. If the output sentences present multiple different answers, carefully determine if the later provided answer is a correction or modification of a previous one. If so, extract this corrected or modified answer as the final response. Conversely, if the output sentences fluctuate between multiple answers without a clear final answer, you should output [No valid answer]. For a question of type "classify_text", if the output sentences do not clearly provide any answer within the specified answer range, then the output should be [No valid answer]. Additionally, if the answer range is a list and the answer given in the output sentences is not included in the answer range, also output [No valid answer]. <?xml version="1.0" ?> <information> <question type="str">{question}</question> <output_sentences type="str">{LLM response}</output_sentences> <key_answer_type type="str">{key answer type}</key_answer_type> <answer_range type="dict"> <type type="str">a list of all possible valid answers</type> <choices type="list">{answer range}</choices> </answer_range> </information> Key extracted answer:
Figure 28: The prompt for automatic annotation by GPT-4: XML form.
E.3 Prompts for Extracting Key Answers Using xFinder
As shown in Figure 29, this is the carefully designed prompt template we used for extracting key answers with xFinder.
You are a help assistant tasked with extracting the precise key answer from given output sentences. You must only provide the extracted key answer without including any additional text. I will provide you with a question, output sentences along with an answer range. The output sentences are the response of the question provided. The answer range could either describe the type of answer expected or list all possible valid answers. Using the information provided, you must accurately and precisely determine and extract the intended key answer from the output sentences. Please don’t have your subjective thoughts about the question. First, you need to determine whether the content of the output sentences is relevant to the given question. If the entire output sentences are unrelated to the question (meaning the output sentences are not addressing the question), then output [No valid answer]. Otherwise, ignore the parts of the output sentences that have no relevance to the question and then extract the key answer that matches the answer range. Below are some special cases you need to be aware of: (1) If the output sentences present multiple different answers, carefully determine if the later provided answer is a correction or modification of a previous one. If so, extract this corrected or modified answer as the final response. Conversely, if the output sentences fluctuate between multiple answers without a clear final answer, you should output [No valid answer]. (2) If the answer range is a list and the key answer in the output sentences is not explicitly listed among the candidate options in the answer range, also output [No valid answer]. Question: {question} Output sentences: {LLM response} Key answer type: {key answer type} Answer range: {answer range} Key extracted answer:
Figure 29: The prompt for extracting key answers by xFinder.